OpenStudy (anonymous):

A spherical weather balloon is being inflated. The radius of the balloon is increasing at the rate of 7 cm/s. Express the surface area of the balloon as a function of time t (in seconds). (Let S(0) = 0.) S(t) = cm^2?

OpenStudy (anonymous):

dr/dt = 7 gives r = 7t+c, c unknown constant s(r) = 4.pi.r^2 so s(t) = 4.pi.(7t+c)^2 sub in given s(0) = 0, to find c, done.

OpenStudy (anonymous):

nope

OpenStudy (campbell_st):

so the answer isn't \[SA = 196 \pi t^2\]

OpenStudy (anonymous):

yep thats it

OpenStudy (campbell_st):

kenny... all you needed to do was some simple algebra... its a worry if you are doing related rates...